# ABScribe: Rapid Exploration of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models

### v1

@@ Acknowledgements.
- McCutchen et al. (2009) Deborah McCutchen, Paul Teske, and Catherine Bankston. 2009. Writing and cognition: Implications of the cognitive architecture for learning to write and writing to learn. In Handbook of research on writing. Routledge, 554–578.
- McPhee (2017) John McPhee. 2017. Draft No. 4: On the writing process. Farrar, Straus and Giroux.
- Murphy et al. ([n. d.]) April Murphy, Stephen E Fancsali, and Steven Ritter. [n. d.]. UpGrade: Scalable Digital Experimentation in Authentic Learning Settings. ([n. d.]).
@@
- Purdy (2014) James P Purdy. 2014. What can design thinking offer writing studies? College Composition and Communication (2014), 612–641.
@@
- Reza et al. (2021) Mohi Reza, Juho Kim, Ananya Bhattacharjee, Anna N. Rafferty, and Joseph Jay Williams. 2021. The MOOClet Framework: Unifying Experimentation, Dynamic Improvement, and Personalization in Online Courses. In Proceedings of the Eighth ACM Conference on Learning @ Scale (Virtual Event, Germany) (L@S ’21). Association for Computing Machinery, New York, NY, USA, 15–26. https://doi.org/10.1145/3430895.3460128
- Rittel (1967) Horst Rittel. 1967. Wicked problems. Management Science,(December 1967) 4, 14 (1967).
@@
- Shakeri et al. (2021) Hanieh Shakeri, Carman Neustaedter, and Steve DiPaola. 2021. Saga: Collaborative storytelling with gpt-3. In Companion Publication of the 2021 Conference on Computer Supported Cooperative Work and Social Computing. 163–166.
- Sharples (2002) Mike Sharples. 2002. How we write: Writing as creative design. Routledge.
@@
- Touvron et al. ([n. d.]) Hugo Touvron, Louis Martin, and Kevin Stone. [n. d.]. Llama 2: Open Foundation and Fine-Tuned Chat Models. ([n. d.]).
- Wu et al. (2023) Yi Wu, Nan Jiang, Hung Viet Pham, Thibaud Lutellier, Jordan Davis, Lin Tan, Petr Babkin, and Sameena Shah. 2023. How Effective Are Neural Networks for Fixing Security Vulnerabilities. In Proceedings of the 32nd ACM SIGSOFT International Symposium on Software Testing and Analysis (Seattle, WA, USA) (ISSTA 2023). Association for Computing Machinery, New York, NY, USA, 1282–1294. https://doi.org/10.1145/3597926.3598135
@@
- Zinsser (2006) William Zinsser. 2006. On writing well: The classic guide to writing nonfiction. New York, NY (2006).

### v3

@@ Acknowledgements.
- McCutchen et al. (2009) Deborah McCutchen, Paul Teske, and Catherine Bankston. 2009. Writing and cognition: Implications of the cognitive architecture for learning to write and writing to learn. In Handbook of research on writing. Routledge, Milton Park, Abingdon, Oxfordshire, United Kingdom, 554–578.
- McPhee (2017) John McPhee. 2017. Draft No. 4: On the writing process. Farrar, Straus and Giroux, 18 West 18th Street New York, NY US.
- Meyer et al. (2022) Selina Meyer, David Elsweiler, Bernd Ludwig, Marcos Fernandez-Pichel, and David E Losada. 2022. Do we still need human assessors? prompt-based gpt-3 user simulation in conversational ai. In Proceedings of the 4th Conference on Conversational User Interfaces. Association for Computing Machinery, New York, NY, USA, 1–6.
- Michel (2020) Holland Michel. 2020. A.: The black box, unlocked: predictability and understandability in military AI. Ginebra, United Nations Institute for Disarmament Research.
- Microsoft (2013) Microsoft. 2013. Reinventing search with a new AI-powered Microsoft Bing and Edge, your copilot for the web. https://blogs.microsoft.com/blog/2023/02/07/reinventing-search-with-a-new-ai-powered-microsoft-bing-and-edge-your-copilot-for-the-web/.
- Min et al. (2023) Bonan Min, Hayley Ross, Elior Sulem, Amir Pouran Ben Veyseh, Thien Huu Nguyen, Oscar Sainz, Eneko Agirre, Ilana Heintz, and Dan Roth. 2023. Recent advances in natural language processing via large pre-trained language models: A survey. Comput. Surveys 56, 2 (2023), 1–40.
- Mirowski et al. (2023) Piotr Mirowski, Kory W. Mathewson, Jaylen Pittman, and Richard Evans. 2023. Co-Writing Screenplays and Theatre Scripts with Language Models: Evaluation by Industry Professionals. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 355, 34 pages. https://doi.org/10.1145/3544548.3581225
- OpenAI (2022) OpenAI. 2022. Introducing ChatGPT. https://openai.com/blog/chatgpt.
@@
- Oppenlaender (2022) Jonas Oppenlaender. 2022. The Creativity of Text-to-Image Generation. In Proceedings of the 25th International Academic Mindtrek Conference (Tampere, Finland) (Academic Mindtrek ’22). Association for Computing Machinery, New York, NY, USA, 192–202. https://doi.org/10.1145/3569219.3569352
- O’Sullivan (2009) Bryan O’Sullivan. 2009. Making sense of revision-control systems. Commun. ACM 52, 9 (2009), 56–62.
@@
- Pichai and Hassabis (2023) Sundar Pichai and Demis Hassabis. 2023. Introducing Gemini: Our largest and most capable AI model. https://blog.google/technology/ai/google-gemini-ai/#sundar-note
- Plaice and Wadge (1993) John Plaice and William W Wadge. 1993. A new approach to version control. IEEE transactions on Software Engineering 19, 3 (1993), 268–276.
- Purdy (2014) James P. Purdy. 2014. What Can Design Thinking Offer Writing Studies? College Composition and Communication 65, 4 (2014), 612–641. http://www.jstor.org/stable/43490875
- Radford et al. (2018) Alec Radford, Karthik Narasimhan, Tim Salimans, Ilya Sutskever, et al. 2018. Improving language understanding by generative pre-training.
@@
- Rittel (1967) Horst Rittel. 1967. Wicked problems. Management Science,(December 1967) 4, 14 (1967), B141–B142.
- Salvagno et al. (2023) Michele Salvagno, Fabio Silvio Taccone, and Alberto Giovanni Gerli. 2023. Artificial intelligence hallucinations. Critical Care 27, 1 (2023), 1–2.
- Sefelin et al. (2003) Reinhard Sefelin, Manfred Tscheligi, and Verena Giller. 2003. Paper prototyping - what is it good for? a comparison of paper- and computer-based low-fidelity prototyping. In CHI ’03 Extended Abstracts on Human Factors in Computing Systems (Ft. Lauderdale, Florida, USA) (CHI EA ’03). Association for Computing Machinery, New York, NY, USA, 778–779. https://doi.org/10.1145/765891.765986
@@
- Shakeri et al. (2021) Hanieh Shakeri, Carman Neustaedter, and Steve DiPaola. 2021. Saga: Collaborative storytelling with gpt-3. In Companion Publication of the 2021 Conference on Computer Supported Cooperative Work and Social Computing. Association for Computing Machinery, New York, NY, USA, 163–166.
- Sharples (2002) Mike Sharples. 2002. How we write: Writing as creative design. Routledge, Milton Park, Abingdon, Oxfordshire, United Kingdom.
- Singh et al. (2023) Nikhil Singh, Guillermo Bernal, Daria Savchenko, and Elena L Glassman. 2023. Where to hide a stolen elephant: Leaps in creative writing with multimodal machine intelligence. ACM Transactions on Computer-Human Interaction 30, 5 (2023), 1–57.
- Smith et al. (2023) Andrew L Smith, Felix Greaves, and Trishan Panch. 2023. Hallucination or Confabulation? Neuroanatomy as metaphor in Large Language Models. PLOS Digital Health 2, 11 (2023), e0000388.
- Snyder (2003) Carolyn Snyder. 2003. Paper Prototyping: The Fast and Easy Way to Design and Refine User Interfaces. Morgan Kaufmann Publishers Inc., San Francisco, CA, USA.
@@
- Tichy (1982) Walter F. Tichy. 1982. Design, implementation, and evaluation of a Revision Control System. In Proceedings of the 6th International Conference on Software Engineering (Tokyo, Japan) (ICSE ’82). IEEE Computer Society Press, Washington, DC, USA, 58–67.
@@
- Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023. Llama 2: Open foundation and fine-tuned chat models.
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Ł ukasz Kaiser, and Illia Polosukhin. 2017. Attention is All you Need. In Advances in Neural Information Processing Systems, I. Guyon, U. Von Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Eds.), Vol. 30. Curran Associates, Inc., Long Beach, CA. https://proceedings.neurips.cc/paper_files/paper/2017/file/3f5ee243547dee91fbd053c1c4a845aa-Paper.pdf
- Walker et al. (2002) Miriam Walker, Leila Takayama, and James A. Landay. 2002. High-Fidelity or Low-Fidelity, Paper or Computer? Choosing Attributes when Testing Web Prototypes. Proceedings of the Human Factors and Ergonomics Society Annual Meeting 46, 5 (2002), 661–665. https://doi.org/10.1177/154193120204600513 arXiv:https://doi.org/10.1177/154193120204600513
- Wang et al. (2023) Yunlong Wang, Shuyuan Shen, and Brian Y Lim. 2023. RePrompt: Automatic Prompt Editing to Refine AI-Generative Art Towards Precise Expressions. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Singapore, 1–29. https://doi.org/10.1145/3544548.3581402
- Wu et al. (2022) Tongshuang Wu, Ellen Jiang, Aaron Donsbach, Jeff Gray, Alejandra Molina, Michael Terry, and Carrie J Cai. 2022. Promptchainer: Chaining large language model prompts through visual programming. In CHI Conference on Human Factors in Computing Systems Extended Abstracts. Association for Computing Machinery, New York, NY, USA, 1–10.
- Yampolskiy (2019) Roman V Yampolskiy. 2019. Unpredictability of AI.
- Yampolskiy (2020) Roman V Yampolskiy. 2020. Unpredictability of AI: On the impossibility of accurately predicting all actions of a smarter agent. Journal of Artificial Intelligence and Consciousness 7, 01 (2020), 109–118.
@@
- Zhang et al. (2023) Lei Zhang, Ashutosh Agrawal, Steve Oney, and Anhong Guo. 2023. VRGit: A Version Control System for Collaborative Content Creation in Virtual Reality. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 36, 14 pages. https://doi.org/10.1145/3544548.3581136
- Zhou et al. (2022) Yongchao Zhou, Andrei Ioan Muresanu, Ziwen Han, Keiran Paster, Silviu Pitis, Harris Chan, and Jimmy Ba. 2022. Large language models are human-level prompt engineers.
- Zinsser (2001) William Knowlton Zinsser. 2001. On writing well: The classic guide to writing nonfiction. Quill/A Harper Collins Books, 195 Broadway New York, NY, USA.
Appendix A Baseline Interface
The Baseline interface featured a conversational AI assistant similar to ChatGPT and the ability to insert AI generated text directly into the document. Refer to Figure 9 for a screenshot.
Appendix B Scenario Descriptions and Prompts
The two task scenarios were described to the participants as follows: